Jo has 37coins( all nickels, dimes and quaters) worth $5.50. She has 4 more quatersthan nickels. How many dimes does Jo have?
Answer by stanbon(75887)

Jo has 37coins( all nickels, dimes and quaters) worth $5.50. She has 4 more quarters than nickels. How many dimes does Jo have?
-----------------
Let number of nickels be "x" ; value of these is 5x cents
Number of quarters is "x+4" ; value of these is 25(x+4)=25x+100 cents
Number of dimes = 37-x-(x+4)=33-2x; value is 10(33-2x)=330-20x cents
--------------------
EQUATION:
value + value + value = 550 cents
5x + 25x+100 + 330-20x = 550
10x = 120
x = 12 (number of nickels)
x+4 = 16 (number of quarters)
37-12-16=9 (number of dimes)
==================
Cheers,
Stan H.